OpenStudy (mendicant_bias):

I'm trying to find the Laplace Transform of a convolution, posted below in a moment.

OpenStudy (mendicant_bias):

I'm looking for \[\mathcal{L}\left\{\vphantom{}t \int\limits_{0}^{t}\sin(\tau)d \tau\right\}\]

OpenStudy (mendicant_bias):

I can't figure out what I'm doing wrong, but I immediately do not understand what to do with the t outside of the integral. @satellite73 , could you help possibly on this?

OpenStudy (mendicant_bias):

The steps I'm taking are this: \[\mathcal{L}\left\{\vphantom{}\int\limits_{0}^{t}\sin(\tau)d \tau\right\}=\mathcal{L}\left\{1\right\}\mathcal{L}\left\{\sin(\tau)\right\}=\frac{1}{s}.\frac{1}{s^2+1}\]

OpenStudy (mendicant_bias):

Hey, @wio . Any chance you could help me with this?

ganeshie8 (ganeshie8):

looks like convolution ?

OpenStudy (mendicant_bias):

Yeah, it does, but I'm not sure how to do it, that's what I'm asking for help on.

OpenStudy (mendicant_bias):

The answer is supposed to be:\[\frac{3s^2+1}{s^2(s^2+1)^2}\]

OpenStudy (mendicant_bias):

What I was doing was separating the integrand into the product of two laplace transforms (by the convolution theorem)-but I must be doing that wrong, because my laplace transforms are correct; the separation into two functions F(s) and G(s) must be wrong.

OpenStudy (anonymous):

If you say that \(t\) is constant with respect to \(\tau\), then:\[ t\int_0^t\sin\tau\;d\tau = \int_0^t t\sin\tau\;d\tau \]

OpenStudy (mendicant_bias):

\[F(t)=\sin(t); \ \ \ G(t) = \ ?\]\[f(\tau)=\sin(\tau); \ g(t - \tau) = \ ?\]

OpenStudy (mendicant_bias):

Yeah, t can be factored in or out, not really sure what to do with it; what is advantageous about factoring t into the integrand?

OpenStudy (anonymous):

t isnt constant tho...its a variable we r gonna sub in...

OpenStudy (anonymous):

treating it like a constant might screw stuff up..

OpenStudy (mendicant_bias):

I'm guessing maybe manipulating the integrand in some manner to get a second function in the from f(t-tau)?

OpenStudy (anonymous):

We sub it into \(\tau\).

OpenStudy (mendicant_bias):

^

OpenStudy (anonymous):

I factored it in because I think \(g(t-\tau)=t\).

OpenStudy (mendicant_bias):

I think you're right, one sec!

OpenStudy (mendicant_bias):

\[\mathcal{L}\left\{\vphantom{}\int\limits_{0}^{t}t \sin(\tau)d \tau\right\}=\mathcal{L}\left\{t\right\}\mathcal{L}\left\{\sin(\tau)\right\}\]

OpenStudy (mendicant_bias):

Whoops, sin(t).

OpenStudy (anonymous):

@hahd \[ t\int_0^tf'(x)\;dx = t(f(t)-f(0)) = tf(t)-tf(0) = \int_0^ttf'(x)\;dx \]

OpenStudy (mendicant_bias):

\[=\frac{1}{s^2}.\frac{1}{s^2+1}\]

OpenStudy (mendicant_bias):

No dice? Really close to the answer, I'm not sure where I'm doing something wrong.

OpenStudy (anonymous):

Wait, maybe since:\[ \frac{d}{dt}tf(t) = tf'(t) +f(t) \]it would be: \[ \int_0^txf'(x)+f(x)\;dx \]

OpenStudy (anonymous):

No, that doesn't work either.

OpenStudy (anonymous):

^ i was thinking this

OpenStudy (anonymous):

leibniz rule..

OpenStudy (anonymous):

\[ \int_0^txf'(x)+f(x)\;dx = tf(t) - 0f(0) = tf(t) \]

OpenStudy (anonymous):

this is a dum question but is tau like 2pi? or its a variable

OpenStudy (anonymous):

lol im really interested now too lol

OpenStudy (anonymous):

It's a variable of integration

OpenStudy (anonymous):

\[ 0<\tau<t \]

OpenStudy (anonymous):

lkk

OpenStudy (anonymous):

how do you know that tho?

OpenStudy (anonymous):

Limits of integration...

OpenStudy (anonymous):

oh gawd kk i get it...

OpenStudy (mendicant_bias):

It's just a dummy variable, it's just there to prevent ourselves from confusing up t in the limits and in the integrand.

OpenStudy (anonymous):

kk but why doesnt leibniz work here

OpenStudy (mendicant_bias):

Well, I don't even understand where you guys are taking a derivative from in the first place.

OpenStudy (anonymous):

its given in the form of an integral of a derivative ??? isnt that the natural thing to do...

OpenStudy (anonymous):

purpose of the derivative in this case is so I can express the antiderivative.

OpenStudy (mendicant_bias):

The laplace transform is an integral operation, and the argument of the Laplace transform is *also* an integral, I don't see where differentiation is coming in at all other than the fact that "d tau" happens to be there, but yeah. I dunno, I pretty much never look at integrals in the manner we are right now, unfamiliar to me.

OpenStudy (mendicant_bias):

@dan815

OpenStudy (mendicant_bias):

(Brb in 5 minutes, time on this PC is running out and need to go to the next door campus building to use another PC)

OpenStudy (zarkon):

\[\mathcal{L}\left\{t\int\limits_{0}^{t}\sin(\tau)d\tau\right\}\] \[\mathcal{L}\left\{t\sin(t)+\int\limits_{0}^{t}\sin(\tau)d\tau\right\}=s\mathcal{L}\left\{t\int\limits_{0}^{t}\sin(\tau)d\tau\right\}\] \[\frac{1}{s}\left[\mathcal{L}\left\{t\sin(t)\right\}+\mathcal{L}\left\{\int\limits_{0}^{t}\sin(\tau)d\tau\right\}\right]=\mathcal{L}\left\{t\int\limits_{0}^{t}\sin(\tau)d\tau\right\}\] \[\frac{1}{s}\left[\frac{2s}{(s^2+1)^2}+\mathcal{L}\left\{\int\limits_{0}^{t}\sin(\tau)d\tau\right\}\right]=\mathcal{L}\left\{t\int\limits_{0}^{t}\sin(\tau)d\tau\right\}\] \[\frac{1}{s}\left[\frac{2s}{(s^2+1)^2}+\frac{\mathcal{L}\left\{\sin(t)\right\}}{s}\right]=\mathcal{L}\left\{t\int\limits_{0}^{t}\sin(\tau)d\tau\right\}\] \[\frac{1}{s}\left[\frac{2s}{(s^2+1)^2}+\frac{\frac{1}{s^2+1}}{s}\right]=\mathcal{L}\left\{t\int\limits_{0}^{t}\sin(\tau)d\tau\right\}\] \[\frac{1}{s}\left[\frac{2s^2}{s(s^2+1)^2}+\frac{s^2+1}{s(s^2+1)^2}\right]=\mathcal{L}\left\{t\int\limits_{0}^{t}\sin(\tau)d\tau\right\}\] \[\frac{3s^2+1}{s^2(s^2+1)^2}=\mathcal{L}\left\{t\int\limits_{0}^{t}\sin(\tau)d\tau\right\}\]

OpenStudy (mendicant_bias):

Alright, I'm back.

OpenStudy (mendicant_bias):

I'm confused about where the first term on the second LHS from the top came from; I'm confused about what operation was done there at all from the original problem.

OpenStudy (mendicant_bias):

@Zarkon (PS, thank you); Just wondering about that first LHS.

OpenStudy (mendicant_bias):

@ganeshie8 , any clue how he did this?

OpenStudy (mendicant_bias):

e.g. where did\[\mathcal{L}\left\{\vphantom{}t \sin(t)+ \int\limits_{0}^{t}\sin(\tau)d \tau\right\}\] come from?

OpenStudy (zarkon):

\[\mathcal{L}\left\{f'(t)\right\}=s\mathcal{L}\left\{f(t)\right\}-f(0)\]

OpenStudy (mendicant_bias):

Alright, cool. Thank you. Going to take a look at this for a little longer.

OpenStudy (mendicant_bias):

\[s \mathcal{L}\left\{\vphantom{}t \int\limits_{0}^{t}\sin(\tau)d \tau\right\}=s\mathcal{L}\left\{f(t)\right\}\]

OpenStudy (mendicant_bias):

Product rule is applied;

OpenStudy (mendicant_bias):

\[f(t)=t \int\limits_{0}^{t}\sin(\tau)d \tau; f'(t)=t \sin(t)+ \int\limits_{0}^{t}\sin(\tau)d \tau\]

OpenStudy (mendicant_bias):

Alright, cool, I got it. Thanks so much, man, that was a really unique/strange approach to apply the definition of the laplace of a derivative.

OpenStudy (zarkon):

even easier it you evaluate the integral \[\mathcal{L}\left\{t\int\limits_{0}^{t}\sin(\tau)d\tau\right\}=\mathcal{L}\left\{-t\cos(t)+t\right\}=\cdots\]

OpenStudy (mendicant_bias):

\[\mathcal{L}\left\{-t \cos(t)+t\right\}=\frac{1}{s^2}-(-1)^{1}\frac{d}{ds}\bigg(\frac{s}{s^2+1}\bigg)\]

ganeshie8 (ganeshie8):

\(*\) : convolution

ganeshie8 (ganeshie8):

i don't see how to use convolution for the remaining integral hmm

ganeshie8 (ganeshie8):

** \[\begin{align}\mathcal{L}\left\{t\int\limits_{0}^{t}\sin(\tau)d\tau\right\}&=\mathcal{L} \left( (t-\tau+\tau)\int\limits_{0}^{t}\sin(\tau)d\tau \right)\\~\\ &=\mathcal{L} \left(\int\limits_{0}^{t} (t-\tau+\tau)\sin(\tau)d\tau \right) \\~\\ &=\mathcal{L} \left(t*\sin(t) \right)+ \mathcal{L}\left(\int\limits_{0}^{t} \tau\sin(\tau)d\tau \right) \\~\\ \end{align}\]

OpenStudy (dan815):

=]

OpenStudy (mendicant_bias):

\[\frac{1}{s^2}-\frac{s^2+1-(2s^2)}{(s^2+1)^2}\]\[\frac{1}{s^2}-\frac{1-s^2}{(s^2+1)^2}\]

OpenStudy (mendicant_bias):

Then just finding an LCD and combining like terms as appropriate:\[\frac{(s^2+1)^2}{s^2(s^2+1)^2}-\frac{(s^2)(1-s^2)}{(s^2+1)^2s^2}\]

OpenStudy (mendicant_bias):

(Lol how about no)

OpenStudy (mendicant_bias):

But seriously, I made an algebraic error somewhere, because it looks like those....yeah, it looks like those...just wait one minute.

OpenStudy (mendicant_bias):

nvm, right hand term is positive altogether.

ganeshie8 (ganeshie8):

why are you not using wolfram to do the donkey work

OpenStudy (mendicant_bias):

\[\frac{s^4+2s^2+1+s^2-s^4}{s^2(s^2+1)^2}=\frac{3s^2+1}{s^2(s^2+1)^2}\]

OpenStudy (mendicant_bias):

Cuz' need to be good at the donkey work for my test in six hours..... :'C

OpenStudy (mendicant_bias):

I'm pretty much having to rapid-fire since last week Monday, constantly do every problem in every section that this test covers prior to 7 PM tonight.

OpenStudy (sidsiddhartha):

man u guys are overdoing it its so easy and u are making it real big :(

OpenStudy (sidsiddhartha):

i can do it in two steps

OpenStudy (sidsiddhartha):

wanna see

ganeshie8 (ganeshie8):

sid can you do it in two steps without evaluating the integral ?

OpenStudy (sidsiddhartha):

ok lets see

OpenStudy (sidsiddhartha):

)

OpenStudy (sidsiddhartha):

we need only two basic properties

OpenStudy (sidsiddhartha):

yes i can do it without any integration @ganeshie8

OpenStudy (mendicant_bias):

:O

ganeshie8 (ganeshie8):

\[\mathcal{L} \left(\int_0^tu\sin u du\right)\] Okie, i have been trying the same and got stuck at above step :O

OpenStudy (sidsiddhartha):

first property: \[L[\int\limits_{0}^{t}F(\tau)d \tau]=\frac{ 1 }{ s }F(s)\\\] next property:\[L[t^n.f(t)]=(-1)^n.\frac{ d^n }{ ds^n }[F(s)]\]

OpenStudy (mendicant_bias):

Siddhartha about to be all like: https://www.youtube.com/watch?v=JGw8DWctAts

OpenStudy (mendicant_bias):

(For now I'm closing this question and moving on since I'm running out of time, heh, but feel free to keep going if you guys want. I'm in crunch-time. Thank you guys so much, again.)

OpenStudy (sidsiddhartha):

so according to those\[L[t.\int\limits_{0}^{t}\sin \tau d \tau]=(-1)^1\frac{ d^1 }{ ds^1 }[\frac{ 1 }{ s }.\frac{ 1 }{ s^2+1 }]=(-1).\frac{ d }{ ds }\frac{ 1 }{ s(s^2+1) }\\=(-1)\frac{ -3s^2-1 }{ s^2(s^2+1)^2 }=\frac{ 3s^2+1 }{ s^2(s^2+1)^2 }\] ta da!!!!!!!!!! done without any integration

ganeshie8 (ganeshie8):

Nice :) still convolution is not used

OpenStudy (sidsiddhartha):

but the question did'nt asked to use convolution it asked to find the Laplace Transform of a convolution

OpenStudy (mendicant_bias):

(Taking a look at this now as well, the best way I think to get a grasp for how to use all techniques is to see how a single problem can be solved in multiple ways)

ganeshie8 (ganeshie8):

Oh right lol i misread the question and messing wid convolution :/

OpenStudy (sidsiddhartha):

and convolution cannot be used in TIME DOMAIN it is used in FREQUENCY(S) DOMIAN right?

OpenStudy (mendicant_bias):

That was pretty clever, the combination of those two properties.

OpenStudy (sidsiddhartha):

we can find inverse laplace using convolution but laplace using convolution is something complicated

OpenStudy (dan815):

hahaha that theme song for sidd

OpenStudy (sidsiddhartha):

lol which one dan?

OpenStudy (dan815):

his youtube video up there

OpenStudy (sidsiddhartha):

lol now i see he looks like a super saiyan :P

OpenStudy (dan815):

how come they make a big deal out of this property 1/s * F(s) lol/ ... like f(t)*g(t)=f(s)g(s)=int(f(tau)(G(t-tau)) tau, its easy to see if u got f(tau) only then g(t) must be a constant or 1, then g(s)= 1/s